OpenStudy (anonymous):

If y = 9x − 7, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(0, −7), (1, 2), (−1, −16)} {(−7, 0), (2, 1), (−16, −1)} {(1, 9), (2, 7), (3, 16)} {(7, 9), (8, 10), (9, 11)}

OpenStudy (anonymous):

?

OpenStudy (anonymous):

could you explain how to find the answer? my brain's gone wacko lol

OpenStudy (anonymous):

I think you have to substitute you multiple choice in to your equation...

OpenStudy (anonymous):

how do you do that?

Theoceangirl (the_ocean_girl):

^ That would probably be the best way to solve. Substitute in the values for x and y, whichever set works for all 3 would be the correct answer

OpenStudy (anonymous):

what the_ocean_girl said

OpenStudy (anonymous):

hmm... so that goes back to what i said before, how do I do that? where do I plug the numbers in?

Theoceangirl (the_ocean_girl):

Let's do the first one together! :) Okay so x would be the input and y is the output. So for A, (0, -7), we would plug 0 in for x and -7 in for y, if this statement turns out to be true, we can move on and check the next group in A. If all 3 groups of numbers fit, we have the right answer. Are you understanding what I'm saying?

OpenStudy (anonymous):

ohh ok thanks!!! so ill figure this out and ill run it by you later :)

Theoceangirl (the_ocean_girl):

No problem! Take your time and let me know what you come up with! :)

OpenStudy (anonymous):

ok thanks!!! thanks @Lovelarap

OpenStudy (anonymous):

@the_ocean_girl

OpenStudy (anonymous):

i havent been able to find an answer ugh and ive tried and tried

OpenStudy (anonymous):

think it's a

OpenStudy (anonymous):

why? can you explain please??

OpenStudy (anonymous):

ok, {(0, −7), (1, 2), (−1, −16)} substitue to y=9x-7 -7=9(0)-7 2=9(1)-7 -16=9(-1)-7

OpenStudy (anonymous):

so then would A be right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

yaya!!!! thanks lol :D :D :D

OpenStudy (anonymous):

your welcome